OpenStudy (anonymous):

The price of blackberries at a farm is $1.56 per pound. Which equation can be used to determine c, the total price of n pounds of blackberries? c = 1.56n c equals 1.56 over n c = 1.56 + n c = 1.56 − n i will give a medals to who helps

OpenStudy (sleepyjess):

Are these the options? \(c=1.56n\\c=\dfrac{1.56}{n}\\c=1.56+n\\c=1.56-n\)

OpenStudy (anonymous):

hi meagan

OpenStudy (anonymous):

yes

OpenStudy (sleepyjess):

Ok, first, which one do you think?

OpenStudy (anonymous):

c =1.56+n

OpenStudy (sleepyjess):

Not quite, lets say you want 5 pounds of blackberries. Do you know which equation now?

OpenStudy (sleepyjess):

\(c=1.56*5\\c=\dfrac{1.56}{5}\\c=1.56+5\\c=1.56-5\)

OpenStudy (anonymous):

b

OpenStudy (sleepyjess):

No, if it were b then you would pay .31 for 5 pounds of blackberries when it is 1.56 for 1 pound.

OpenStudy (sleepyjess):

I will give you a hint, the answer should be greater than 7.

OpenStudy (anonymous):

so it the last 1

OpenStudy (sleepyjess):

1.56-5 is -3.44

OpenStudy (sleepyjess):

This is a test?

OpenStudy (sleepyjess):

Have you read the Code of Conduct and Terms and Conditions? If so, then you would know that asking questions from tests is not allowed here.

OpenStudy (sleepyjess):

Because asking questions from tests is cheating.